OpenStudy (anonymous):

Simplify the given expression. 4 divided by the quantity of 3 minus 2i four thirteenths sixteen thirteenths the quantity of 12 plus 8i, divided by 13 the quantity of 12 minus 8i, divided by 13

OpenStudy (anonymous):

@Todd_The_Best

OpenStudy (anonymous):

3-2= 1, 4/1 = 4 4/13 16/13 = 1 3/16 20/13 = 1.53 0.307

OpenStudy (anonymous):

Please give me the medal and follow me.

OpenStudy (anonymous):

happy..

OpenStudy (anonymous):

Thanx